OpenStudy (anonymous):

Multiply. 1 and one-third times 5 and five-eighths 5 and five twenty-fourths 7 and one-half 5 and six−elevenths 8

OpenStudy (saifoo.khan):

1 and one-third times 5 and five-eighths \[1\frac13 \times 5 \frac58\]\[\frac43 \times \frac{45}{8}\]Multiply.

OpenStudy (anonymous):

7/2 i think

OpenStudy (saifoo.khan):

15/2

OpenStudy (anonymous):

o thankz

OpenStudy (anonymous):

Multiply. 5 and one-fourth times twelve equals sixty and one-fourth sixty-three sixty-three and one-fourth sixty-one and one-fourth

OpenStudy (saifoo.khan):

try them. :P

OpenStudy (anonymous):

63

OpenStudy (saifoo.khan):

for which one?

OpenStudy (anonymous):

this one Multiply. 5 and one-fourth times twelve equals sixty and one-fourth sixty-three sixty-three and one-fourth sixty-one and one-fourth